# ERASER: Machine Unlearning in MLaaS via an Inference Serving-Aware Approach

### v2

@@ 6.4. Temporal Distribution of Requests
Please refer to Appendix I for additional experimental results. Moreover, Appendix H delves into scenarios where requesters might have malicious intentions.
@@ 7. Conclusion and Future Work
Appendix A What we do not achieve in this paper
We briefly discuss what we do not aim to achieve in this paper.
Regarding extra security and privacy risk: We mainly focus on ensuring that data owners who have submitted unlearning requests do not suffer from extra privacy risk due to the postponed unlearning execution. However, we do not address other sources of privacy threats, such as general privacy attacks considered in MIA literature where any data owners including those has not submitted unlearning request can be the victim (Shokri et al., 2017; Salem et al., 2019; Li and Zhang, 2021; Liu et al., 2022a). We also do not address the issue of attackers impersonating data owners to submit fraudulent unlearning requests, as we consider this to fall under the domain of user authentication, which is orthogonal to the research objectives of this paper.
Appendix B Deferred proof of Theorem 4.2
@@ Appendix A What we do not achieve in this paper
Appendix C Deferred proof of Theorem 5.1
@@
Appendix D Deferred proof of Theorem 5.2
@@
Appendix E Discussions
E.1. Why not Differential Privacy?
To certify inference consistency without actual machine unlearning execution, we need the concept of model robustness with respect to small changes in the training dataset. Roughly speaking, this refers to the property that two ML models trained on neighboring datasets (i.e., two datasets that have most of the data in common except a few) have the same inference results. This naturally evokes the idea of utilizing Differential Privacy (DP) (Dwork et al., 2006) techniques for model training. However, the primary challenge with the DP-based approach is that the robustness offered by DP allows only for approximate unlearning, rather than exact unlearning. As the number of unlearning request keeps increasing during the MLaaS service, the privacy bound of DP may be breached with the growth of to-be-erased data. Approximate unlearning methods with DP-like bounds may face the same challenge. Additionally, the noise introduced by DP can significantly decrease the model’s utility.
E.2. Difference between Inference Consistency and Certified Robustness
Theorem 4.2 draws inspiration from the certified robustness defense against data poisoning attacks (Levine and Feizi, 2021; Wang et al., 2022; Jia et al., 2021). However, as demonstrated in Appendix F, straightforwardly applying the certified robustness from (Levine and Feizi, 2021) results in a worse certification condition. In other words, ERASER would no longer certify on inference samples it should have. In fact, Theorem F.1 in Appendix F can be regarded as having the certification condition $2\gamma_{1}+2\gamma_{2}+2\gamma_{3}\leq\Gamma$, which is apparently more strict than ours $2\gamma_{1}+0\gamma_{2}+1\gamma_{3}\leq\Gamma$. As a result, our certification analysis is tailored to the machine unlearning under MLaaS and achieves a more fine-grained certification condition.
E.3. Discussion of uncertification ratio
Our experimental results indicate that the $p_{uc}$ value is typically less than 0.01, meaning that the waiting time for DIMP does not exceed 1% of SISA, as shown in Figure 5. We can also qualitatively analyze $p_{uc}$ based on the characteristics of the model:
Appendix F Straightforward adaptation from (Levine and Feizi, 2021)
Theorem F.1 ().
@@ Appendix F Straightforward adaptation from (Levine and Feizi, 2021)
Since $\gamma=\gamma_{1}+\gamma_{2}+\gamma_{3}$, it is obvious that $2\gamma\geq 2\gamma_{1}+\gamma_{3}$. Thus, it indicates that Theorem 4.2 provides a more fine-grained certification bound than Theorem F.1, which also shows that our certification analysis is tailored to machine unlearning in MLaaS.
Appendix G Algorithm Frameworks
G.1. The Overall Framework of ERASER
1:Training dataset $\bm{\mathscr{D}}=\{\mathbf{x}_{1},\dots,\mathbf{x}_{N}\}$
2:
5:Launch inference service with $F=\{f_{1},\dots,f_{K}\}$
6:
7:for $t\in[1,2,\dots]$ do
9:         Record it as a pending unlearning request
13:         if satisfy certified inference consistency condition  then
15:         else
16:             Process in accordance with specific design options.
17:         end if
18:    end if
19:end for
20:Responses for inference and unlearning requests.
G.2. Consistency Certification Mechanism
Algorithm 2 present the details of the inference consistency certification mechanism.
Algorithm 2 Inference Consistency Certification Mechanism
2:$certification_{\tt{q}}=True$
3:for $k\in[1,2,\dots,K]$ do
4:    Inference with constituent model: $f_{k}^{\tt O}(\mathbf{z})$
5:    if $\bm{\mathscr{S}}_{k}^{\tt O}\cap\tt{U}_{k}^{t}\neq\varnothing$  then
6:         Mark as $\bm{\mathscr{S}}_{k}^{t}\neq\bm{\mathscr{S}}_{k}^{\tt O}$
7:    end if
8:end for
10:for $\forall y_{b}\in\mathbb{Y},y_{b}\neq y_{a}$ do
15:    if $2\gamma_{1}+\gamma_{3}>\Gamma_{b}$ then
16:         $certification_{\tt{q}}=False$
17:         Break
18:    end if
19:end for
20:if $certification_{\tt{q}}$ then
22:else
23:    The consistency of $\tt{q}$ can not be certified
24:end if
25:$certification_{\tt{q}}$
Appendix H Malicious Requests and Mitigation
In addition to the toy attacks in Sec.3.2, we present two new attacks that can be launched by queriers and data owners through malicious requests, aiming to consume more of the server’s computational resources and slow down its response time. Experimental results are shown in Appendix I.3.
H.1. Hard-to-Classify Inference Requests
@@
H.2. Scattered Unlearnings Requests
@@
Appendix I Additional Experiments
@@ I.1. Model Accuracy
I.2. Evaluation of Parallel Capacity
@@
I.3. Malicious Inference and Unlearning Requests

### v3

@@ 6.4. Temporal Distribution of Requests
Please refer to the appendices in arXiv version for additional experimental results, including the scenarios where requesters might have malicious intentions.
@@ 7. Conclusion and Future Work
Acknowledgements.
This work is supported by the National Key Research and Development Program of China (No. 2022YFE0113200), the National Natural Science Foundation of China (No. 62206207 & No. 62072395), and the Joint Fund of the National Natural Science Foundation of China (No. U20A20178).
Appendix A Deferred proof of Theorem 4.2
@@ Acknowledgements.
Appendix B Deferred proof of Theorem 5.1
@@
Appendix C Deferred proof of Theorem 5.2
@@
Appendix D Straightforward adaptation from (Levine and Feizi, 2021)
Theorem D.1.
@@ Appendix D Straightforward adaptation from (Levine and Feizi, 2021)
Theorem 4.2 draws inspiration from the certified robustness defense against data poisoning attacks (Levine and Feizi, 2021; Wang et al., 2022; Jia et al., 2021). However, straightforwardly applying the certified robustness from (Levine and Feizi, 2021) results in a worse certification condition. In other words, ERASER would no longer certify on inference samples it should have. In fact, Theorem D.1 in Appendix D can be regarded as having the certification condition $2\gamma_{1}+2\gamma_{2}+2\gamma_{3}\leq\Gamma$, which is apparently more strict than ours $2\gamma_{1}+0\gamma_{2}+1\gamma_{3}\leq\Gamma$. As a result, our certification analysis is tailored to the machine unlearning under MLaaS and achieves a more fine-grained certification condition.
Appendix E Malicious Requests and Mitigation
In addition to the toy attacks in Sec.3.2, we present two new attacks that can be launched by queriers and data owners through malicious requests, aiming to consume more of the server’s computational resources and slow down its response time. Experimental results are shown in Appendix F.3.
E.1. Hard-to-Classify Inference Requests
@@
E.2. Scattered Unlearnings Requests
@@
Appendix F Additional Experiments
@@ F.1. Model Accuracy
F.2. Evaluation of Parallel Capacity
@@
F.3. Malicious Inference and Unlearning Requests